### v2

# Field-orientation-dependent magnetic phases in GdRu2Si2 probed with muon-spin spectroscopy

Abstract
Centrosymmetric GdRu2Si2 exhibits a variety of multi-Q magnetic states as a function of temperature and applied magnetic field, including a square skyrmion-lattice phase. The material’s behavior is strongly dependent on the direction of the applied field, with different phase diagrams resulting for fields applied parallel or perpendicular to the crystallographic $c$ axis. Here, we present the results of muon-spin relaxation ($\mu^{+}$) measurements on single crystals of GdRu2Si2. Our analysis is based on the computation of muon stopping sites and consideration of zero-point motion effects, allowing direct comparison with the underlying spin textures in the material. Using transverse-field $\mu^{+}$ with fields applied along either the [001] or [100] crystallographic directions, we distinguish between the magnetic phases in this system via their distinct muon response, providing additional evidence for the skyrmion and meron-lattice phases, while also suggesting the existence of RKKY-driven muon hyperfine coupling. Zero-field $\mu^{+}$ provides clear evidence for a transition between two distinct magnetically-ordered phases at 39 K.
I Introduction
Skyrmions are topologically-protected vortices of magnetization that exist in some magnetically-ordered materials [1]. The search for new materials that host a magnetic skyrmion state has been motivated by the numerous potential applications of skyrmions in the areas of ultra-efficient information storage and spintronics [2]. Skyrmions are typically stabilized by the Dzyaloshinskii–Moriya interaction (DMI), which can only occur in systems that lack inversion symmetry. Magnets based on Gd (possessing 4f7 core states with $S=7/2$, $L=0$) are promising candidates for realizing non-collinear spin structures such as skyrmions, since the moments show non-preferential orientation. Two Gd-based magnets where skyrmions have recently been discovered are Gd2PdSi3 [3] and Gd3Ru4Al12 [4]. In these centrosymmetric materials, skyrmions have been proposed to be stabilized by either short- or long-range magnetic frustration rather than DMI. This can lead to interesting competition with other magnetic phases, and potentially new behavior of the skyrmion state. Skyrmions have also been observed in GdRu2Si2 [5], another centrosymmetric Gd-based magnet. However, the structure of GdRu2Si2, shown in Fig. 1(a), comprises alternating square-lattice Gd and Ru2Si2 layers and therefore, unlike for Gd2PdSi3 and Gd3Ru4Al12, short-range geometrical frustration cannot be used to explain the formation of the skyrmion lattice (SkL). Magnetic resonant x-ray scattering (RXS) and real space using Lorentz transmission electron microscopy reveal the existence of a double-Q magnetic state, corresponding to a square SkL [5]. This is distinct from the hexagonal SkL found in Gd2PdSi3 and GdRu4Al12, and in the vast majority of the noncentrosymmetric skyrmion-hosting materials. Moreover, the existence of a square SkL has only been demonstrated [5] for applied fields $\bm{B}_{0}\parallel$ [001], for which the SkL phase is stabilized at $\bm{B}_{0}\approx$ 2 T. For fields in the $ab$ plane, a different phase diagram is obtained [6, 7]. The magnetic phase diagrams for $\bm{B}_{0}\parallel$ [001] and $\bm{B}_{0}\parallel$ [100] have been studied using RXS [7], with these measurements revealing a rich variety of double-Q magnetic structures, including the antiferroic order of meron/antimeron-like textures. Neutron diffraction measurements on single-crystal and polycrystalline 160GdRu2Si2 found, in addition to scattering at the principal propagation vectors $\bm{q}_{1}$ and $\bm{q}_{2}$, higher-order satellites at $\bm{q}_{1}+2\bm{q}_{2}$, which prompted a revision of the zero-field magnetic ground state to a double-Q constant-moment magnetic structure, with one-dimensional topological charge density [8]. At small values of applied field, an additional phase has been identified just below $T_{\mathrm{N}}$ whose properties have not been reported.
00footnotetext: See Supplemental Material at [URL will be inserted by publisher] for further details of the the density functional theory calculations and simulations of the magnetic field distributions, which includes Refs. [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42].
Here, we use muon-spin spectroscopy ($\mu^{+}$) [9] to probe the magnetic phases in GdRu2Si2 as a function of magnetic field and temperature. The $\mu^{+}$ technique has been shown to be sensitive to the distribution of magnetic fields arising from the SkL phase in Cu2OSeO3 [10], as well as those found in the helical [11] and conical [12] phases of skyrmion-hosting MnSi. Implanted muons have also been shown to effectively probe the emergent dynamics associated with Bloch SkL, such as in Cu2OSeO3 and CoxZnyMn20-x-y [13] and in Gd2PdSi3 [14], or those accompanying a Néel SkL, as found in GaV4S8 and GaV4Se8 [15, 16]. Using transverse-field (TF) $\mu^{+}$ for magnetic fields applied along either [001] or [100], we show here how the response of the implanted muon reflects the occurrence of skyrmion and other double-Q magnetic phases. Zero-field (ZF) $\mu^{+}$ measurements provide clear evidence for a transition between two distinct magnetically-ordered phases at 39 K, showing that the phase reported just below $T_{\mathrm{N}}=45$ K is ordered throughout the bulk of the material. Our modelling also affords an estimate of the low-temperature ordered moment of $4.8(1)\mu_{\mathrm{B}}$ and shows the influence of long-range Ruderman-Kittel-Kasuya-Yosida (RKKY) coupling in the system, via the muon’s coupling to the electronic degrees of freedom.
The paper is structured as follows: in Sec. II we present the results of muon stopping site calculations using density functional theory (DFT), in Sec. III we present the results of $\mu^{+}$ measurements on GdRu2Si2 and discuss these with support from simulations of the magnetic field at the muon site for candidate magnetic structures, and finally we present our conclusions in Sec. IV. Additional details can be found in the Supplemental Material (SM) [17].
II Muon site calculations
Implanted muons are sensitive to the distribution of magnetic fields at the muon sites. Knowledge of these sites can facilitate a quantitative understanding of $\mu^{+}$ measurements. Progress has recently been made in computing stopping sites using density DFT [18], although treatment of quantum-mechanical effects such as zero-point motion (ZPM) has hitherto been limited to special cases [19, 9]. We have used DFT methods to calculate the muon sites in GdRu2Si2 [17] and find two crystallographically-distinct candidate sites [Fig. 1(a)]. The calculated entanglement witnesses [19] at each of these sites places both of these muon states firmly within the weakly-bound adiabatic limit. However, on taking into account the quantum zero-point energy, a muon at site 2 (the higher-energy state of the two) has sufficient energy to overcome the energy barrier between its classical location and that of site 1 [Fig. 1(c)]. Site 2 is therefore not stable under quantum fluctuations of the muon position. While the barrier from site 1 to site 2 is sufficiently high that we do not expect muons at site 1, we also need to consider the possiblity of quantum mechanical hopping. By considering the effective one-dimensional potential for the muon along the minimum-energy path between the classical turning points, i.e., where the black and red lines in Fig. 1(c) intersect, we used the WKB approximation to estimate a tunnelling rate $\approx$0.05 Hz for muons at site 1 [17], which is significantly longer than the timescale over which measurements are made. We therefore conclude that there is a single crystallographically-distinct stable stopping site in this material and we plot the displacements of the nearby ions in the vicinity of this site in Fig. 1(b). The muon does not introduce any significant perturbations to its local environment, with all displacements $<0.1$, and is therefore expected to be a faithful probe of the magnetism in this system. These calculations do not take into account the possible effect of charge density waves, which have been proposed to result from the coupling between itinerant electrons and localized moments in this system [20], and could result in small perturbations to the muon site.
Figure 1: (a) Two crystallographically distinct muon stopping sites in GdRu2Si2, obtained using DFT. (b) Muon-induced displacements of ions as a function of distance from the lowest-energy muon site. (c) Energy barriers between muon sites. Red and blue lines indicate the total (classical plus quantum, zero-point) energy of the muon at sites 1 and 2, respectively.
III Results and Discussion
III.1 TF $\bm{\mu}^{+}$ measurements
Transverse-field $\mu^{+}$ measurements were made on aligned crystals of GdRu2Si2 after zero-field cooling. In these measurements, the muon spin precesses about the sum of the external magnetic field, which is oriented perpendicular to the initial muon-spin direction, and any internal magnetic fields present in the sample. Two sets of measurements were made: one with the crystals oriented such that applied field $\bm{B}_{0}$ pointed along [001] and another set with the crystals oriented such that $\bm{B}_{0}$ pointed along the [100] direction.
We first consider the spectra measured at $T=40$ K, where we can access the field-polarized state. Fourier transform (FT) $\mu^{+}$ spectra measured at different fields are shown in Fig. 2 for each of the orientations. For $\bm{B}_{0}\parallel$ [001] the spectra comprise a peak close to the applied field and a secondary peak at lower fields which becomes more prominent as the field is increased. (We note that the central peak is split in two for all of the TF spectra measured on this material, which could result from a fraction of the muons stopping in parts of the sample where the internal field almost, but not completely, cancels, such as, for example, domain walls or surfaces.) For $\bm{B}_{0}\parallel$ [100] there is an additional peak at fields larger than $B_{0}$ that is absent for the other orientation. We therefore fit the spectra for fields along [001] and [100] to the sum of two or three Lorentzians, respectively, and plot the behavior of the peak centers in Figs. 2(d) and 2(h). We see that the magnitudes of the field shifts at the two peaks initially increases approximately linearly with the applied field, but then begins to level off, suggesting that the magnetization is starting to saturate. We also note the absence of any satellite peaks for $B_{0}=0.5$ T $\parallel$ [001] in Fig. 2(a), with the magnetic phase diagram for this orientation [shown in Fig. 3(i)] having a phase boundary between 0.5 and 1 T at this temperature. This suggests that the low-field, higher-temperature magnetically ordered state gives rise to a broad distribution of magnetic fields at the muon site, relaxing the contribution of muons stopping in the sample out of the TF spectra, a scenario which is consistent with our ZF spectra, which show a large relaxation rate at this temperature. On the other hand, spectra measured at 0.5 and 1.0 T for $\bm{B}_{0}\parallel[100]$, shown in Figs. 2(e) and 2(f), respectively, are expected to occupy the same magnetic phase [as seen in our ac susceptibility measurements shown in Fig. 2(j)] and both show satellite peaks. This suggests the zero-field ordered state survives up to higher fields applied along [001] than for those along [100].
Figure 2: Fourier transform TF $\mu^{+}$ spectra measured at $T=40$ K for applied magnetic fields along (a–c) [001] and (e–g) [100]. Positions of peaks as a function of applied field $B_{0}$ along [001] and [100] are shown in (d) and (h), respectively. Red and blue lines in (h) correspond to fits of the internal field at muon sites lying along the $b$ axis or $a$ axis respectively, as a function of the extracted Gd magnetic moment. (i) Magnetic moment of the Gd ions at 40 K, obtained from the separation of the peaks in the muon spectra (red circles) and magnetometry (black). (j) Real part of the ac susceptibility as a function of temperature and magnetic field for applied magnetic fields along [100]. Circles on the phase diagram correspond to the FT $\mu^{+}$ spectra presented in this paper, with colors corresponding the lines used to the spectra.
The number of peaks observed for each of the field orientations reflects the symmetry of the muon stopping site. The local magnetic field at the muon site is the sum of the externally applied field and dipolar, Lorentz, demagnetizing and hyperfine fields. For Gd moments along $c$, muons at site 1 will all experience an internal field of $(-0.063-0.140N+7.77A)\mu_{\mathrm{Gd}}$ T along $c$, where the first term is sum of the dipolar and Lorentz fields, the second term is the demagnetizing field, with the demagnetizing factor $N$ depending on the sample shape, and the final term is the hyperfine field, which is proportional to the hyperfine coupling constant $A$, here measured in Å-3. All of these contributions to the internal field are proportional to the Gd magnetic moment $\mu_{\mathrm{Gd}}$, which in this expression is measured in Bohr magnetons. On the other hand, for moments along $a$, instances of site 1 lying on the $b$ axis will experience a field $(0.295-0.140N+7.77A)\mu_{\mathrm{Gd}}$ T along $a$, whereas the crystallographically equivalent site lying on the $a$ axis will instead experience a field $(-0.091-0.140N+7.77A)\mu_{\mathrm{Gd}}$ T. For $\bm{B}_{0}\parallel$ [100] the separation of the satellite peaks, $\Delta B$, is therefore related to $\mu_{\mathrm{Gd}}$ by $\Delta B=0.386\mu_{\mathrm{Gd}}$ T. The development of $\mu_{\mathrm{Gd}}$ as a function of applied magnetic field is shown in Fig. 2(i). We have also measured the Gd moment as a function of magnetic field using magnetometry [17] [Fig. 2(i)]. The values of $\mu_{\mathrm{Gd}}$ obtained using these approaches show reasonable agreement, though we note that those obtained from magnetometry are consistently larger and have a slightly different functional dependence with applied field. Taking the sample geometry to be a flat plate with a perpendicular magnetization, we assume $N\approx 1$. Fitting either the positive or negative and field shifts for $\bm{B}_{0}\parallel$ [100] to a straight line in $\mu_{\mathrm{Gd}}(B_{0})$ [blue or red lines in Fig. 2(h), respectively] yields $A=-0.004~{}\mathrm{\AA}^{-3}$. This negative hyperfine contribution can be explained by the spin-polarization of conduction electrons through the RKKY interaction, which has been predicted to play an important role in stabilizing the skyrmion lattice in GdRu2Si2 [21]. We note that while the multiplicity of the peaks for fields along [100] is not specific to the candidate sites found in our calculations, it does rule out sites with axial symmetry, such as the site at (0.5, 0.5, 0) that has been proposed for the isostructural compound CeRu2Si2 [22], which would exhibit only a single peak.
We now turn to the spectra measured at $T=5$ K, where several distinct magnetic structures exist as a function of applied magnetic field [5, 7]. FT $\mu^{+}$ spectra are shown in Fig. 3 for each orientation. For $B<2.25$ T, the spectra comprise a split peak, close to the applied field. For $\bm{B}_{0}\parallel$ [001] [Figs. 3(b-d)], the spectra contain additional spectral weight at fields below the applied field, whereas for $\bm{B}_{0}\parallel$ [100], we see a pair of satellite peaks, above and below $B_{0}$. As before, this reflects the symmetry of the muon site for each field orientation. However, the features in the spectra are much broader for $T=5$ K, suggesting that the system adopts magnetic structures that give rise to a wider distribution of fields. These become sharper at higher applied fields, as the field-polarized component of the magnetic structure becomes more significant.
Figure 3: Fourier transform TF $\mu^{+}$ spectra measured at $T=5$ K for applied magnetic fields along (a–d) [001] and (e–h) [100]. Black lines indicate simulations of the spectra. (i) Magnetic phase diagram for applied fields along [001], using phase boundaries obtained in Ref. [5].
[…]
For the spectra measured at $B_{0}=1.0$ T [Figs. 3(a, e)], we see only a sharp peak at the applied magnetic feature, with no other notable features. This peak is due to muons stopping outside of the sample, suggesting that muons stopping inside the sample experience a broad distribution of magnetic fields and therefore have their contribution to the TF spectra dephased. This is consistent with our simulations based on the constant-moment magnetic structure proposed in Ref. [8] [denoted CM in Fig. 3(i)], which leads to a Fourier spectrum with weight spread across a wide frequency window with no notable peaks. The simulated spectrum for $\bm{B}_{0}\parallel$ [001] does, however, predict some additional spectral weight at fields around 0.5 T above the applied field, which is not seen in experiment, although the overall broadness of the spectrum means that this would likely be difficult to resolve in practice.
We first focus on the spectra measured for $\bm{B}_{0}\parallel$ [001]. For $B_{0}=2.25$ T, GdRu2Si2 has been proposed to host a square SkL phase [5]. The corresponding $\mu^{+}$ spectrum in Fig. 3(b) exhibits two peaks at fields below the applied field, whose locations show good agreement with those simulated for an approximate realization of a square SkL. Our simulation also predicts a peak just above the applied field, suggesting that a fraction of the implanted muons will be at sites where the internal field is well aligned with the external field. Due to the anisotropic nature of the dipolar tensor at the muon site, a magnetic structure with moments polarized along the $c$ axis results in a dipolar field that points in the opposite direction to the magnetization, and hence the external field, as seen in Fig. 2(c). This therefore highlights the role of the incommensurate components of the magnetic moments in giving rise to a local field at the muon site aligned with the externally-applied field. However, we note that this high-field feature is not resolved in the measured spectrum, possibly due to being too close to the central peak.
The 2.25 T spectrum is qualitatively different from those measured at 2.5 [Fig. 3(c)] and 4 T [Fig. 3(d)] for this orientation, allowing us to clearly distinguish the response of the muon to the SkL from that arising from the other multi-Q magnetic structures that are realized in this system. For the higher fields along [001], this has been proposed to be the checkerboard antiferroic order of meron/anti-meron-like textures (MLc) [7]. The $B_{0}\geq 2.5$ T spectra both comprise a satellite peak below the applied field, although the shape of the satellite is distinct in each case. Our simulated spectra based on the MLc magnetic structure show good agreement with experiment for both values of $B_{0}$. The difference between the spectra measured at 2.25 [Fig. 3(b)] and 2.5 T [Fig. 3(c)] can be explained by the fact that the helical components of the SkL magnetic structure rotate into and out of the $c$-axis direction, either aligning or antialigning with the ferromagnetic component, whereas the modulated parts of the moment in the MLc phase remain within the $ab$ plane (and therefore perpendicular to the ferromagnetic component). Thus the former magnetic structure more effectively splits the spectral weight into distinct peaks, whereas the latter only results in broadening of the low-field feature.
Turning to the spectra measured for $\bm{B}_{0}\parallel$ [100], we note that, in contrast to $\bm{B}_{0}\parallel$ [001], applied fields with magnitudes $B_{0}=2.25$ T and $B_{0}=2.5$ T in this orientation are expected to result in a single-Q conical magnetic structure. The spectra measured in this phase [Fig. 3(f) and Fig. 3(g)] comprise pairs of satellite peaks, one above $B_{0}$ and one below. These features are fairly broad and somewhat difficult to resolve in experiment for $B_{0}=2.25$ T, but become sharper at $B_{0}=2.5$ T. Our simulated spectra can be interpreted as a pair of shifted Overhauser distributions [23] resulting from a helical magnetic structure propagating along the $a$ axis, which leads to different magnetic fields for muons sitting along the $a$ axis to those lying on the $b$ axis. An additional ferromagnetic component shifts the centers of these distributions, such that one is centered on a field above $B_{0}$ and one on a field below, and also results in a small splitting of the peaks within this distribution, which is particularly noticeable for the distribution at fields below $B_{0}$. It is difficult to reconcile these spectra with those observed in experiment. The peaks observed in the experimental spectra suggest the presence of a significant field-polarized ferromagnetic component, but these are broadened, rather than experiencing the splitting that our simulations predict. We therefore suggest that these spectra are incompatible with the proposed single-Q magnetic structure. It is possible that the magnetic structure for this phase needs to be revised, and further work on this phase would be instructive.
At $B_{0}=4$ T, an applied field parallel to [100] gives rise to another checkerboard phase, where the modulations in magnetic moment are rotated relative to those for $\bm{B}_{0}\parallel$ [001], in order to remain perpendicular to the applied field [7]. For both field orientations, the magnetic structures at $B_{0}=4$ T contain substantial field-polarized components, so the locations and multiplicity of the peaks mostly reflect the symmetry of the muon sites, as was discussed for the spectra measured at $T=40$ K.
Figure 4: Zero-field $\mu^{+}$ spectra measured at (a) $1.8$ K, (b) $38$ K, (c) $43$ K and (d) $55$ K. Temperature dependence of (e) the muon procession frequency $f$ and the relaxation rates (f) $\lambda_{1}$ and (g) $\lambda_{2}$ in Eq. (1).
III.2 ZF $\bm{\mu}^{+}$ measurements
Zero-field $\mu^{+}$ measurements were made on two aligned GdRu2Si2 single crystals. The ZF $\mu^{+}$ spectra exhibit a very quickly relaxing oscillating asymmetry. The first 0.1 $\mu$s of the spectra were fitted to the function
$A(t)=A_{1}\exp(-\lambda_{1}t)\cos(2\pi ft+\phi)+A_{2}\exp(-\lambda_{2}t),$ (1)
where the oscillating and non-oscillating components of the asymmetry have amplitudes $A_{1}$ and $A_{2}$ and relaxation rates $\lambda_{1}$ and $\lambda_{2}$, respectively. The amplitudes $A_{1}$ and $A_{2}$ and the phase $\phi$ were globally refined, with values $A_{1}=12.6(3)\,\%$, $A_{2}=5.56(4)\,\%$ and $\phi=-17(2)^{\circ}$. The temperature dependence of the variable parameters is shown in Fig. 4. The muon precession frequency $f$ serves as a magnetic order parameter for the system and exhibits a discontinuity in its derivative that suggests a transition between two distinct magnetic phases, both ordered throughout the bulk. The frequencies $f(T)$ were fitted to
$f(T)=\begin{cases}f_{1}[1-(T/T_{\mathrm{c}})^{\alpha}]^{\beta_{1}}&\text{for }0\leq T\leq T_{\mathrm{N}1}\\ f_{2}(1-T/T_{\mathrm{N}2})^{\beta_{2}}&\text{for }T_{\mathrm{N}1}\leq T\leq T_{\mathrm{N}2}\\ \end{cases},$ (2)
where $f_{2}=\frac{[1-(T_{\mathrm{N}1}/T_{\mathrm{c}})^{\alpha}]^{\beta_{1}}}{(1-T_{\mathrm{N}1}/T_{\mathrm{N}2})^{\beta_{2}}}$, ensuring that $f(T)$ is continuous at $T=T_{\mathrm{N}1}$. This fitting yields $f_{1}=129(3)$ MHz, $\alpha=4.06(8)$, $\beta_{1}=0.49(8)$, $\beta_{2}=0.36(2)$, $T_{\mathrm{c}}=39.5(1)$ K, and transition temperatures $T_{\mathrm{N}1}=38.8(1)$ K and $T_{\mathrm{N}2}=44.95(1)$ K. The critical exponent $\beta_{2}=0.36(2)$ for the transitions between the higher-temperature ordered phase and the paramagnetic phase shows good agreement with the value $\beta=0.366$ expected for a 3D Heisenberg magnet [24]. The transition at $T_{\mathrm{N}1}$ can also be seen in the relaxation rates, with $\lambda_{1}$ [Fig. 4(f)] and $\lambda_{2}$ [Fig. 4(g)] exhibiting discontinuous increases and decreases, respectively.
[…]
Figure 5: Fits to a model based on the constant-moment solution for data measured at (a) 1.8 K, (b) 25 K and (c) 38 K. The blue line in (c) indicates a fit to an alternative model based on a sine magnetic structure. (d) Gd magnetic moments and (e) fluctuation rates $\nu$ extracted using simulations of the $\mu^{+}$ spectra as a function of temperature.
The extracted fluctuation rates $\nu\approx$3 MHz are significantly smaller than the relaxation rates $\lambda_{1}$ that accompany the relaxing part of the asymmetry, indicating that this relaxation is mostly due to the distribution of static magnetic fields resulting from the incommensurate magnetic structure. The discontinuous behavior of $\lambda_{1}$ at $T_{\mathrm{N1}}$ is therefore evidence for a significant change in the magnetic structure at this temperature. The fluctuation rates $\nu$ obtained from our simulations are similar in magnitude to the slow relaxation rate $\lambda_{2}$, consistent with the interpretation that this relaxation is due to dynamics flipping the muon spin at sites where the local magnetic field is parallel to the muon spin, in the slow fluctuation limit. We note that the relative magnitudes of $A_{1}$ and $A_{2}$ are consistent with those of the oscillating and non-oscillating parts, respectively, of the polarization function obtained by our model.
Unlike the magnetic ground state [8], the higher-temperature magnetically ordered phase has not been unambiguously identified. However, a recent powder neutron powder diffraction study [25] proposed three candidate single-q magnetic structures for this phase, with these measurements being unable to distinguish between these. We therefore simulated $\mu^{+}$ spectra based on each of the models and fit these to the data measured at $T\geq 38$ K. All of these models provided similar quality fits to the data, with a so-called sine magnetic structure with moments along the [011] direction being the best, and also being the only model that gives a magnetic moment that decreases monotonically with temperature, as shown in Fig. 5(d). (We note that for these sine magnetic structures, which do not have a constant magnetic moment amplitude, $\mu_{\mathrm{Gd}}$ represents the maximum value of the magnetic moment.) This magnetic structure also provides a better fit to the spectrum measured at $T=38$ K than the constant-moment solution, as indicated by the blue line in Fig. 5(c). Our modelling also indicates that the fluctuation rate increases on approaching $T_{\mathrm{N2}}$ from below, as shown in Fig. 5(e), but remains within the slow fluctuation limit. This suggests that the distribution of static magnetic fields resulting from an incommensurate magnetic structure is also responsible for the majority of the relaxation observed in this higher-temperature ordered phase.
IV Conclusion
We have illustrated how implanted muons are sensitive to the magnetic phases in GdRu2Si2, providing evidence for skyrmion and meron-lattice magnetic phases. In contrast, $\mu^{+}$ spectra measured for fields along [100] are inconsistent with the previously proposed single-Q magnetic structure, suggesting that this phase needs to be revised. We demonstrate a significant, negative hyperfine contribution to the local magnetic field at the muon site, due to conduction electrons becoming spin-polarized through the RKKY interaction. Zero field measurements as a function of temperature demonstrate a transition between two distinct bulk magnetically ordered states at $T_{\mathrm{N}1}=38.8(1)$ K.
Data presented in this paper will be made available via Ref. [26].
Acknowledgements.
Part of the work was carried out at the Swiss Muon Source, Paul Scherrer Institute, Switzerland and we are grateful for the provision of beamtime and experimental support. We acknowledge computing resources provided by Durham Hamilton HPC. This work was funded by the EPSRC UK Skyrmion Project under Grant No. EP/N032128/1. The work at the University of Warwick was further supported by EPSRC grant EP/T005963/1. Work at Oxford was funded by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee (Grant No. EP/X025861/1). AHM is grateful to STFC and EPSRC for the provision of a studentship under Grant No. EP/T518001. MG acknowledges the financial support of the Slovenian Research and Innovation Agency through Program No. P1-0125 and Projects No. J1-50008, J1-50012, and and N1-0356.